A thermodynamic process follows this relation: P = kV⁻². If the initial pressure is P₁ and the final pressure is P₂, then what is the work done on the gas? Heat can flow through the container.


1
Expert's answer
2022-01-31T07:34:52-0500

As we know, work is


"W=nRT\ln\dfrac{V_1}{V_2}=nRT\ln\sqrt\dfrac{P_2}{P_1}."


Pay attention that we had to express pressure from the equation


"P=kV^{-2}."